Fermi National Laboratory


Accelerator Update for July 27, 2005 - July 29, 2005

Wednesday July 27
The day shift began with Operations monitoring store 4301 and a small Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, NuMI, and Meson MCenter and MTest experiments taking beam.

At 8:37 AM, Operations reported a trigger problem with the Main Injector (MI) 52 kicker scope. An expert investigated and performed some tests, but didnít find a problem.

Operations put the Recycler into shot setup at 1:52 PM to stash antiprotons. The stash ended and Pbar resumed stacking at 2:51 PM.

MCenter and MTest resumed taking beam by 3:01 PM.

MTest requested to have their beam turned off at 9:32 PM.

Thursday July 28
The midnight shift began with Operations monitoring store 4301 and the stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

At 2:04 AM, the Tevatron (TeV) suffered a multi-house quench. Later, it was suspected that a kicker prefire had caused the quench.

A TeV expert suggested that the separators be put into a conditioning mode. Operators did this at 3:22 AM.

At 5:52 AM, Operations reported that the Pbar emittances were growing. An Operator notified an expert. Stacking was halted.

Operators began the TeV turn on at 6:23 AM.

Pbar resumed stacking at 6:53 AM. The cause for the emittance blowup was never found.

The TeV checkout was completed by 7:19 AM.

Operations put the TeV through a dry squeeze at 7:56 AM.

Operations put the TeV into shot setup at 9:07 AM.

Operations established combination store 4304 at 12:02 PM with an initial luminosity of 102.2E30. Pbar resumed stacking.

MCenter resumed taking beam at 12:50 PM.

At 2:48 PM, store 4304 was lost due to studier error.

Operators prepared CDF and D0 for an access into their collision halls at 3:13 PM.

CDF came out of access at 4:15 PM.

D0 came out of access at 6:24 PM.

TeV experts began a study period at 7:33 PM. They ended their study at 10:35 PM.

Operations put the TeV into shot setup at 11:48 PM.

Friday July 29
The midnight shift began with the TeV in shot setup, and with MiniBooNE and NuMI taking beam.

Operations established store 4308 at 2:23 AM with an initial luminosity of 67.43E30. Pbar resumed stacking.

MCenter resumed taking beam at 2:27 AM.

At 7:25 AM, Operations reported that a Booster RF station (BRF19) had trip off twice. It reset okay.

The Plans for Friday and the Weekend
The plans for today and the weekend are to stack, stash, and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 7/29/2005   email Fermilab